OpenStudy (anonymous):

The amount of money in an investment is modeled by the function A(t)=750(1.0258)^t. The variable A represents the investment balance in dollars, and t the number of years. What is the doubling time for the investment?

OpenStudy (wolf1728):

We started with $750.00 so we want to know when A = 1.500.00 1,500 = 750 * (1.0258)^t 1,500 / 750 = 1.0258^t Taking logs of both sides: log (2) = t * log(1.0258) t = log (2) / log(1.0258) t = .301029996 / 0.0110626947 t = 27.2112720289 years And to check that answer, try this page: http://www.1728.org/compint.htm